a pool can be filled water using pipe a in 4hrs drained using pipe b in 10 hours. if pipe b is accidentally left open while the pool is being filled how, long will it take to fill the pool

Answer by mananth(16946)

a pool can be filled water using pipe a in 4hrs drained using pipe b in 10 hours. if pipe b is accidentally left open while the pool is being filled how, long will it take to fill the pool
pipe A ----- 4 hours
so it does 1/4 of job in 1 hour,
Pipe B ----------- 10 hours
so it does 1/10 of the job in 1 hour.
..
1/4 -1/10 is the job done in 1 hour when both are open
..
(10-4)/40= 6/40 = 3/20 of the job
..
so it will take 20/3 hours to fill the pool.
=6.66 hours